OpenStudy (anonymous):

PLEASE HELP, I WILL GIVE BEST RESPONSE, AND MEDAL! The formula for the area of the triangle is A=1/2 bh. where b is the base and h is the height. what is the are of the triangle shown here 16cm and 18cm triangle

OpenStudy (anonymous):

the area is 144. 16*18/2=144

OpenStudy (anonymous):

So, we have the formula A=1/2bh.. Plug in the numbers. Now multiply them 16x18=288 Now divide 288 in half 288/2 = 144 So, as you can see, the area in the triangle is 144 cm.

OpenStudy (dn_20):

We have the formula A = 1/2bh Plug in the number A = 1/2 (16) (18) Divide A = 1/2 (288) So A = 144 So the area of the triangle is 144